### v2

@@ Abstract
Purpose While existing approaches excel at recognising current surgical phases, they provide limited foresight and intraoperative guidance into future procedural steps. Similarly, current anticipation methods are constrained to predicting short-term and singular events, neglecting the dense and sequential nature of surgical workflows. To address these needs and limitations, we propose SWAG (Surgical Workflow Anticipative Generation), a framework to combine phase recognition and anticipation, using a generative approach for surgical workflow guidance.
Methods This paper investigates two distinct decoding methods—single-pass (SP) and auto-regressive (AR)—to generate sequences of future surgical phases at minute intervals over long horizons of up to 60 minutes. We propose a novel embedding approach using prior knowledge to enhance the accuracy of phase anticipation. Additionally, our anticipative framework offers remaining time regression and proposes a regression-to-classification (R2C) method. SWAG’s performance was evaluated on the Cholec80 and AutoLaparo21 datasets.
Results Our single-pass model with prior knowledge embeddings (SP*) achieves 49.8% mean accuracy over 18-minute anticipation on AutoLaparo21, while the simple SP with R2C extension reaches 56.6% mean accuracy over the same horizon on Cholec80. Moreover, our approach outperforms existing methods on the phase remaining time regression task, achieving weighted mean absolute errors of 0.32 and 0.48 minutes for 2- and 3-minute horizons, respectively.
Conclusion SWAG demonstrates versatility across classification and regression tasks and creates a temporal continuity between surgical workflow recognition and anticipation. While further studies are required to understand the impact of generative-based anticipation intraoperatively, our method provides steps towards this direction. Project: https://maxboels.github.io/swag.
@@ 1 Introduction
Current approaches predominantly focus on surgical phase recognition, identifying the present phase, step, or action [3, 4, 5]. These methods are valuable for postoperative analysis but offer limited assistance in real-time intraoperative decision-making. They cannot anticipate future events to allow for dynamic planning adjustments, which could potentially improve surgical outcomes. Remaining Surgery Duration (RSD) estimation has been actively researched for over two decades [6, 7, 8, 9, 10, 11], demonstrating clinical interest in long-term predictions—i.e. predicting the remaining time until the end of the surgery. Other studies explore surgical workflow anticipation, such as predicting the next phases and instrument occurrences [12, 13]. However, these methods are limited by their requirement to predict a single event per class, disregarding scenarios where multiple future occurrences may exist within a fixed time horizon.
Generative models, such as GPT models [14], address this challenge by generating sequences of tokens that can vary in both length and class frequency. In surgical workflow anticipation, auto-regressive (AR) [14] and single-pass (SP) decoding models [15] demonstrate the potential to predict continuous, coherent sequences of surgical actions. Unlike conventional remaining time regression approaches, generative models output multiple tokens, extending the observed sequence of events to a plausible future representation of the surgical workflow.
@@
    We introduce SWAG, a generative model that combines surgical phase recognition and anticipation for long-term and dense future prediction of workflows.
@@
    We provide an extensive comparison of two generative decoding methods—single-pass and auto-regressive—alongside two tasks—classification and regression—to generate sequences of future surgical phases.
@@
    We propose a novel prior knowledge embedding mechanism using class-conditional probabilities to initialise the input tokens, enhancing anticipation accuracy.
@@
    We conduct a comprehensive evaluation on the Cholec80 and AutoLaparo21 datasets, demonstrating SWAG’s performances on different surgical procedures.
@@ 3.1 Task Formulation
Our primary task is to predict future surgical phases over a horizon of $H=60$ minutes, conditioned on observed frames $X_{t}=\{x_{0},\ldots,x_{t}\}$ where $x_{i}\in\mathcal{X}_{T}$ (all video frames). We define the mapping function as follows:
$f_{\theta}:x_{\leq t}\mapsto\bigl{(}y_{t+n_{1}\cdot 60},\ldots,y_{t+n_{H}\cdot 60}\bigr{)}$ (1)
where $x_{\leq t}$ denotes the sequence of frames up to time $t$. The anticipation indices $n_{i}$ range from $1$ to $H$, corresponding to the number of minutes into the future for which we predict the surgical phase $y_{t+n_{i}\cdot 60}$.
For regression, we predict the remaining time until each phase’s next occurrence within $[0,H]$ minutes, where 00 indicates the phase is currently active, and $H$ means it will not occur within the horizon. Our setup follows [12, 13]. We predict phases every 60 seconds based on our ablations (see Fig. 2 in SM).
Figure 1: Overview of the SWAG-SP* model architecture. The model processes surgical video data in two main paths: (1) A recognition path with $L=24*60$ frame embeddings and $M=24$ compressed frame representations are encoded into $D=512$-dimensional embeddings through a Vision Encoder, a Windowed Self-Attention (WSA) encoder and a Compression and max-Pooling (CP) bottleneck to recognise the current surgical phase with a classification head (Lossrecognition). (2) A generative path samples $H$ times from future class probabilities extracted from the training set. These probabilities are combined with temporal position embeddings to form input tokens $Q_{t}=\{{q_{1},q_{2},...,q_{H}}\}$, and passed to a Decoder that conditions the compressed frame embeddings i.e. context tokens $K_{t}=\{{k_{1},k_{2},...,k_{M}}\}$ to predict $H$ future surgical phases at $60$ seconds intervals (Lossanticipation). The timeline shows the model operating on past-present observed frames from phases P1 (orange) and P2 (green) and predicting probabilities $H\times C$ of future phases.
@@ 3.2 SWAG Model Architecture
Our proposed model, illustrated in Figure 1, comprises a Vision Encoder followed by a Windowed Self-Attention encoder (WSA), then a linear Compression and max-Pooling mechanisms (CP) [5], and finally a future Decoder module that can operate either in a single-pass (SP) or an auto-regressive (AR) approach. While our model uses classification for the phase recognition task, it can use classification or regression for phase anticipation. This architecture also includes a novel Prior Knowledge Embedding relying on the predicted class $\hat{y_{0}}$ at anticipation index 00 and the index position of the anticipated time horizons $\mathcal{N}_{H}$.
@@
Window Self-Attention (WSA). To recognise the observed frame at time $t$, we use an input $F_{t}={\{f_{t-l},\ldots,f_{t}}\}$ of length $l=1440$s or 24 minutes and latent representations with $d=768$ dimensions. We then use a short sliding window of width $w=20$s to perform self-attention over the input sequence with no overlap and get an output sequence of the same length and dimensionality, $E_{t}={\{e_{t-l},\ldots,e_{t}}\}$.
Compression and Pooling (CP). We evaluate two temporal pooling methods i.e. global key-pooling [5] and our interval-pooling with single-pass and auto-regressive decoders (see Fig. 3 in SM). We select global key-pooling [5] as the best of both methods. This approach compresses the $E_{t}={\{e_{t-l},\ldots,e_{t}}\}$ features into a lower dimensional latent space $d=32$ and aggregates long-term information using max-pooling into the last $M=24$ frame embeddings. A linear layer returns the compressed features to their original dimensionality $D=512$ as $K_{t}={\{k_{1},\ldots,k_{M}}\}$. This bottleneck helps extract global information and passes it to local features [5]. Alternatively, we tried interval-pooling that does max-pooling over 60-second intervals to preserve the positional information. This method has a fixed start time at $t-l$, while the right-side index is extended by $60$ seconds for each new token, effectively compressing $60$ frame embeddings into a single token, yielding $M$ context tokens. $M$ corresponds to the full length divided by 60-second intervals, therefore compressing 24 minutes into 24 tokens, i.e. $\frac{1440}{60}$.
@@
1) SWAG-SP (Single-Pass): The decoder receives $H$ input tokens $Q_{t}=\{{q_{1},...,q_{H}}\}$ and generates $H$ output tokens $S_{t}=\{{s_{1},...,s_{H}}\}$ in a single forward pass, each token represents $60$ seconds over $H$ minutes. During training, we use a special teacher-forcing approach, using the ground-truth current phase $y_{t}$ to sample from the correct prior distributions $P(y_{t+n_{i}\cdot 60}=j\mid y_{t}=i)$ at future minute index $n_{i}$, $H$ times. The $Q_{t}$ inputs are initialized with prior knowledge corresponding to conditional future class probabilities extracted from the training set (see Sec. 1.2 in SM) and sinusoidal positional encoding. This generative approach is conditioned on the $M$ context tokens $K_{t}={\{k_{1},\ldots,k_{M}}\}$ once to generate all future tokens $S_{t}$, allowing for efficient parallel computing and reducing inference time.
@@
2) SWAG-AR (Auto-Regressive): Our decoder architecture builds upon the GPT-2 framework [14] which takes a single input sequence $K_{t}$, and predicts the next token with masked self-attention. Teacher forcing is not used as our approach uses the $M$ learned frame embeddings from the recognition channel rather than embedding ground-truth labels into inputs. SWAG-AR uses $M$ context tokens as input, with predictions for $k_{t+1}$ conditioned on the spatiotemporal context tokens $K_{t}=\{k_{1},\ldots,k_{M}\}$. This context token decoding method iteratively expands during inference until generating all $H$ future tokens.
@@
Recognition Head (Rec. Head). We first fuse the key-pooled features $K_{t}$ with the windowed self-attention features $E_{t}=\{{e_{t-M},...,e_{t}}\}$ using a skip connection [5]. The resulting fused feature vectors are then passed through a classification layer to predict the future $M$ class probabilities for $\{\hat{p_{1}},\ldots,\hat{p_{M}}\}$. The final prediction $\hat{y_{0}}$ for frame $x_{t}$, takes the last frame’s class probabilities, $\hat{p_{t}}$.
@@
Anticipation Head (Ant. Head). The future phase classification task predicts outputs with shape $H\times C$, with $\{\hat{p_{1}},\ldots,\hat{p_{H}}\}$. The final linear layer (LN) $\hat{y_{0}}$ for frame $x_{t}$, takes the last frame’s class probabilities, $\hat{p_{t}}$.
@@
Note that, the regression task outputs $1\times C$ values for the remaining time until the next phases, including the end-of-surgery (eos) class. While remaining time regression could be performed via auto-regressive decoding with a single iteration, we opt for a single-pass decoder since a single token can represent specific transitions over time, and also due to the stronger performance on the classification task. Nevertheless, future work could explore generating multiple transitions per class (e.g., one for each auto-regressively generated token) to narrow potential performance gaps relative to an SP scheme.
@@
SWAG-SP* (with Prior Knowledge Embedding). Our extracted prior knowledge captures the conditional probabilities of phase transitions, derived exclusively from the training set. Specifically, we compute the probability of transitioning from the current class $i$ at time $t$ to future classes $j$ at different anticipation times $t+n_{i}\cdot 60$, denoted as $P(y_{t+n_{i}\cdot 60}=j\mid y_{t}=i)$. These probabilities form a tensor $P$, where each entry $P_{i,j,n_{i}\cdot 60}$ represents the likelihood of transitioning between classes at each anticipation time. We use these distributions to initialise future token embeddings, assigning each future token at $t+n_{i}\cdot 60$ the probability vector $P_{i,:,n_{i}\cdot 60}$. While these priors are specific to each surgical procedure type, they could potentially be transferred between similar procedures, which we plan to investigate in future work.
@@
SWAG-SP-R2C (Regression-to-Classification). We demonstrate a possible mapping from regression outputs to classification via our regression-to-classification (R2C) method, which converts continuous time predictions into discrete phase intervals by sorting and binning predicted times. This approach can be used if one wishes to derive a phase sequence from a model originally trained for regression. Specifically, remaining time predictions are mapped to a discrete sequence of $H$ bins containing the future $H$ predicted class integers $Y_{t}=\{\hat{y_{1}},\ldots,\hat{y_{H}}\}$, in ascending order of occurrence. We use the current predicted class $\hat{y_{0}}$ to fill the sequence up to $\hat{y_{1}}$.
We use cross-entropy and mean squared error losses for the classification and regression tasks, respectively.
@@
Figure 2: Single-Pass regression and classification tasks. The regression task (top) uses a single $1\times D$ input token to predict the remaining time before each next class occurrence, which can be transformed into a classification task using our Regression-to-Classification (R2C) method. The direct classification task (bottom) uses $H$ input tokens to predict $H$ minutes in the future.

### v3

@@ Abstract
Purpose While existing approaches excel at recognising current surgical phases, they provide limited foresight and intraoperative guidance into future procedural steps. Similarly, current anticipation methods are constrained to predicting short-term and single events, neglecting the dense, repetitive, and long sequential nature of surgical workflows. To address these needs and limitations, we propose SWAG (Surgical Workflow Anticipative Generation), a framework that combines phase recognition and anticipation using a generative approach.
Methods This paper investigates two distinct decoding methods—single-pass (SP) and auto-regressive (AR)—to generate sequences of future surgical phases at minute intervals over long horizons. We propose a novel embedding approach using class transition probabilities to enhance the accuracy of phase anticipation. Additionally, we propose a generative framework using remaining time regression to classification (R2C). SWAG was evaluated on two publicly available datasets, Cholec80 and AutoLaparo21.
Results Our single-pass model with class transition probability embeddings (SP*) achieves 32.1% and 41.3% F1 scores over 20 and 30 minutes on Cholec80 and AutoLaparo21, respectively. Moreover, our approach competes with existing methods on phase remaining time regression, achieving weighted mean absolute errors of 0.32 and 0.48 minutes for 2- and 3-minute horizons.
Conclusion SWAG demonstrates versatility across generative decoding frameworks and classification and regression tasks to create temporal continuity between surgical workflow recognition and anticipation. Our method provides steps towards intraoperative surgical workflow generation for anticipation. Project: https://maxboels.github.io/swag.
@@ 1 Introduction
Current approaches predominantly focus on surgical phase recognition, identifying the present phase, step, or action [3, 4, 5]. These methods are valuable for postoperative analysis but offer limited assistance in real-time intraoperative decision-making. They cannot anticipate future events to allow for dynamic planning adjustments, which could potentially improve surgical outcomes. Remaining Surgery Duration (RSD) estimation has been actively researched for over two decades [6, 7, 8, 9, 10, 11], demonstrating clinical interest in long-term predictions—i.e. predicting the remaining time until the end of the surgery. Other studies explored surgical workflow anticipation, such as predicting the next phases and instrument occurrences [12, 13]. However, these methods are limited by their requirement to predict a single event per class, disregarding scenarios where multiple future occurrences may exist within a fixed time horizon.
Generative models, such as GPT models [14], address this challenge by generating sequences of tokens that can vary in both length and class frequency. In surgical workflow anticipation, auto-regressive (AR) [14] and single-pass (SP) decoding models [15] demonstrate the potential to predict continuous sequences of surgical actions. Unlike conventional remaining time regression approaches, generative models output multiple tokens, extending the observed sequence of events to a plausible future representation of the surgical workflow.
@@
    We introduce SWAG, a generative model that combines surgical phase recognition and anticipation, for long-term sequential future prediction of the workflow.
@@
    We provide an extensive comparison of two generative decoding methods—single-pass and auto-regressive—alongside two tasks—classification and regression—to generate continuous sequences of present and future surgical phases.
@@
    We propose a novel prior knowledge embedding method using future class transition probabilities improving predictive performances.
@@
    We conduct a comprehensive evaluation on the Cholec80 and AutoLaparo21 datasets, demonstrating SWAG’s capabilities on different surgical procedures.
@@ 3.1 Task Formulation
Our primary task is to predict the current and future surgical phases over a horizon of $N=h_{N}$ minutes, conditioned on observed frames $X_{t}=\{x_{0},\ldots,x_{t}\}$, where $x_{i}\in\mathcal{X}_{T}$ represents all video frames from the surgical procedure. We define the mapping function as follows:
$f_{\theta}:X_{\leq t}\mapsto\bigl{(}y_{t+h_{0}\cdot 60},\ldots,y_{t+h_{N}\cdot 60}\bigr{)}$ (1)
where $X_{\leq t}$ denotes the sequence of observed frames up to time $t$. The anticipation sequence $\{h_{n}\}_{n=1}^{N}$ is a predefined set of time steps, where each $h_{n}$ corresponds to the number of minutes into the future for which we predict the surgical phase. The sequence is defined as:
$h_{0}=0,\quad h_{1},h_{2},\dots,h_{N}\in\{1,2,\dots,N\}$ (2)
where $h_{0}$ represents the current phase prediction and $h_{N}$ defines the maximum anticipation horizon.
For the time regression task, we predict the remaining time until each phase’s next occurrence within a fixed horizon of $[0,N]$ minutes, where 00 indicates the phase is currently active, and $N$ means it will not occur within the horizon. Our setup follows [12, 13]. We predict phases every 60 seconds based on our ablations (see Fig. 4 in SM).
Figure 1: Overview of the SWAG-SP* model architecture. The model processes surgical video data in two main paths: (1) A recognition path with $L=24*60$ frame embeddings and $M=24$ compressed frame representations are encoded into $D=512$-dimensional embeddings through a Vision Encoder, a Windowed Self-Attention (WSA) encoder and a Compression and max-Pooling (CP) bottleneck to recognise the current surgical phase with a classification head (Lossrecognition). (2) A generative path samples $N$ times from future class probabilities extracted from the training set. These probabilities are combined with temporal position embeddings to form input tokens $Q_{t}=\{{q_{1},q_{2},...,q_{N}}\}$, and passed to a Decoder that conditions the compressed frame embeddings i.e. context tokens $K_{t}=\{{k_{1},k_{2},...,k_{M}}\}$ to predict $N$ future surgical phases at $60$ seconds intervals (Lossanticipation). The timeline shows the model operating on past-present observed frames from phases P1 (orange) and P2 (green) and predicting probabilities $N\times C$ of future phases.
@@ 3.2 SWAG Model Architecture
Our proposed model, illustrated in Figure 1, includes a Vision Encoder followed by a Windowed Self-Attention encoder (WSA), then a linear Compression and max-Pooling block (CP) [5], and finally a future Decoder module that can operate either in a single-pass (SP) or an auto-regressive (AR) approach. While our model uses classification for the phase recognition task, it can use classification or regression for phase anticipation. This architecture also includes a novel Prior Knowledge Embedding relying on the predicted current class $\hat{y}_{t+h_{0}\cdot 60}$ and future indices from $h_{1}$ to $h_{N}$.
@@
Window Self-Attention (WSA). To recognise the observed frame at time $t$, we use the extracted features as input sequence $F_{t}={\{f_{t-L},\ldots,f_{t}}\}$ with a length of $L=1440$ and $D=768$ dimensions. We then use a sliding window of width $W=20$ to perform self-attention over the extracted image features with no overlap between windows and get an output sequence of the same length and dimensionality, $E_{t}={\{e_{t-L},\ldots,e_{t}}\}$.
Compression and Pooling (CP). We implement two temporal pooling strategies: global key-pooling [5] and interval-pooling, which are used differently in our single-pass and auto-regressive decoders. The single-pass method exclusively uses global key-pooling, while the auto-regressive approach leverages both methods. Specifically, the recognition branch in both models uses global key-pooling. For global key-pooling, the temporal features $E_{t}={e_{t-L},\ldots,e_{t}}$ are first projected into a lower-dimensional latent space ($d=32$) as $E^{\prime}_{t}$ using a linear layer, then compressed via a cumulative max-pooling approach into $M$ tokens $K_{t}={k_{1},\ldots,k_{M}}$. Following [5], this involves first computing a single pooled representation of the compressed features: $p=\max{\{e^{\prime}_{t-L},\ldots,e^{\prime}_{t-M}\}}$ Then, for each of the most recent $M$ frames, we compute a cumulative maximum: $k_{m}=\max{\{p,e^{\prime}_{t-M+1},\ldots,e^{\prime}_{t-M+m}\}}$ for $m=1,2,\ldots,M$ This creates tokens that progressively incorporate more recent information while maintaining context from all previous frames. The auto-regressive approach, trained with causal masking on input sequences, requires temporal consistency between the training sequence and the generated outputs tokens during inference. Therefore, we developed interval-pooling to aggregate spatial information at regular temporal intervals. Interval-pooling performs max-pooling over consecutive 60-second intervals, with a fixed start time at $t-L$. For each new token, the right-side index extends by $60$ seconds, effectively compressing $60$ frame embeddings into a single token. This yields $M$ context tokens, where $M$ corresponds to the sequence length divided by the 60-second interval (e.g., compressing 24 minutes into 24 tokens: 1440 seconds60 $\frac{1440\text{ seconds}}{60\text{ seconds}}$). This approach ensures a progressive and cumulative aggregation of information which is consistent with the time interval in the generated sequence during inference.
@@
1) SWAG-SP (Single-Pass): The vanilla transformer decoder receives $N$ input tokens $Q_{t}=\{{q_{1},...,q_{N}}\}$ and generates $N$ output tokens $S_{t}=\{{s_{1},...,s_{N}}\}$ in a single forward pass, each token represents $60$ seconds over $N$ minutes. During training, we use a special teacher-forcing approach, using the ground-truth current phase $y_{t}$ to sample from the correct conditional probability distribution $P(y_{t+h_{n}\cdot 60}=j\mid y_{t}=i)$ at future minute index $h_{n}$, $N$ times. The $Q_{t}$ inputs are initialized with prior knowledge corresponding to conditional future class probabilities extracted from the training set (see Sec. 1.2 in SM) and sinusoidal positional encoding. This generative approach is conditioned using cross-attention between the context tokens $K_{t}={\{k_{1},\ldots,k_{M}}\}$ and the $Q_{t}$ inputs to generate all future tokens $S_{t}$ in a single forward pass, allowing for efficient parallel computing and reducing inference time.
@@
2) SWAG-AR (Auto-Regressive): We use the GPT-2 model as our auto-regressive decoder [14] which uses causal masking on the features extracted with interval-pooling to predict the next token in the sequence. Teacher forcing is not used as our approach uses the learned frame embeddings from the recognition task rather than embedding ground-truth labels as inputs. During inference, the decoder iteratively uses past predictions as inputs until its generate $N=h_{N}$ future tokens as its anticipated horizon.
@@
Recognition Head (Rec. Head). We fuse the key-pooled features $K_{t}$ with the windowed self-attention features $E_{t}=\{{e_{t-M-1},...,e_{t}}\}$ using a skip connection [5]. The resulting fused feature vectors are then passed through a classification layer to predict $M$ class probabilities $\{\hat{p_{1}},\ldots,\hat{p_{M}}\}$, which estimate the class labels assigned to the input frames $\{{x_{t-M-1},\ldots,x_{t}}\}$.
@@
Anticipation Head (Ant. Head). We use a linear layer (LN) and softmax to predict future phase probabilities $\{\hat{p_{1}},\ldots,\hat{p_{N}}\}$ with shape $N\times C$.
@@
Note that, the regression task outputs $1\times C$ values for the remaining time until the next phases, including the End-Of-Surgery (EOS) class. While remaining time regression could be performed via auto-regressive decoding with a single iteration, we opt for a single-pass decoder since a single token can represent specific transitions over time, and also due to the stronger performance on the classification task. Nevertheless, future work could explore generating multiple transitions per class (e.g., one for each generated token) to narrow potential performance gaps relative to an SP scheme.
@@
SWAG-SP* (with Prior Knowledge Embedding). Our extracted prior knowledge captures the conditional probabilities of phase transitions at specific future times and is exclusively derived from the training set. Specifically, we compute the probability of having a future class $j$ at time step $t+h_{n}\cdot 60$ given the current observed class $i$, denoted as $P(y_{t+h_{n}\cdot 60}=j\mid y_{t}=i)$. These probabilities form a tensor $P$, where each entry $P_{i,j,h_{n}\cdot 60}$ represents the likelihood of transitioning between classes at each anticipated time. We use these probabilities to initialise future token embeddings, assigning each future token at $t+h_{n}\cdot 60$ the probability vector $P_{i,:,h_{n}\cdot 60}$.
@@
SWAG-SP-R2C (Regression-to-Classification). We demonstrate a possible mapping from regression outputs to classification via our regression-to-classification (R2C) method, which converts continuous time predictions into discrete phase intervals by sorting and binning predicted times. This approach can be used if one wishes to derive a phase sequence from a model originally trained for regression. Specifically, remaining time predictions for each class are mapped to a discrete sequence of $N$ bins containing the predicted class integers $\hat{Y}_{t}=\{\hat{y}_{0},\hat{y}_{1},\ldots,\hat{y}_{N}\}$, in ascending order of occurrence. We use the current predicted class $\hat{y}_{0}$ to fill the sequence up to the first next class occurrence and use cross-entropy and mean squared error losses for the classification and regression tasks, respectively.
@@
Figure 2: Single-Pass regression and classification tasks. The regression task (top) uses a single $1\times D$ input token to predict the remaining time before each next class occurrence, which can be transformed into a classification task using our regression-to-classification (R2C) method. The direct classification task (bottom) uses $N$ input tokens to predict $N$ minutes in the future.